# AverageTime: Enhance Long-Term Time Series Forecasting with Simple Averaging

### v3

@@ Abstract
Long-term time series forecasting focuses on leveraging historical data to predict future trends. The core challenge lies in effectively modeling dependencies both within sequences and channels. Convolutional Neural Networks and Linear models often excel in sequence modeling but frequently fall short in capturing complex channel dependencies. In contrast, Transformer-based models, with their attention mechanisms applied to both sequences and channels, have demonstrated strong predictive performance. Our research proposes a new approach for capturing sequence and channel dependencies: AverageTime, an exceptionally simple yet effective structure. By employing mixed channel embedding and averaging operations, AverageTime separately captures correlations for sequences and channels through channel mapping and result averaging. In addition, we integrate clustering methods to further accelerate the model’s training process. Experiments on real-world datasets demonstrate that AverageTime surpasses state-of-the-art models in predictive performance while maintaining efficiency comparable to lightweight linear models. This provides a new and effective framework for modeling long time series.
Machine Learning, ICML
@@ 1 Introduction
Long-term time series prediction involves forecasting future trends over extended periods based on historical changes. This approach is crucial in various fields such as weather forecasting (Ma et al., 2023), traffic prediction (Long et al., 2024), and power demand estimation (Baruque et al., 2019). The extended forecast horizon and the complex correlations across sequences and channels pose significant modeling challenges. Traditional methods often fall short of capturing sequence and inter-channel relationships. Consequently, the primary methodologies in this field have shifted to deep learning models.
@@
The core issue in long-term time series analysis is extracting dependencies within sequences and channels, which are crucial for improving model performance and robustness in multi-channel prediction. Various methods have been developed to capture these correlations from time series data. Widely employed techniques include Transformers (Wu et al., 2021; Zhang & Yan, 2022; Zhou et al., 2022, 2021; Nie et al., 2023; Liu et al., 2024; Wang et al., 2024b), which apply attention mechanisms to capture correlations both within sequences and across channels; Convolutional Neural Networks (CNN) (Wang et al., 2022; Wu et al., 2023), which use 1D or multidimensional convolutions to capture these dependencies; and structures based on Multilayer Perceptrons(Zeng et al., 2023; Chen et al., 2023; Lin et al., 2024b, a; Wang et al., 2024a; Li et al., 2023), such as DLinear, which decompose sequences and apply multiple linear layers to predict future series. Among these, the iTransformer (Liu et al., 2024), which captures channel dependencies using channel embeddings and an attention structure, has achieved significant performance improvements on benchmark datasets, especially in datasets with a large number of channels. However, as the importance of channel correlations has been increasingly recognized (Zhao & Shen, 2024), many models have not been able to effectively capture these correlations, leading to a decline in model performance. This inadequacy becomes particularly evident in scenarios with a large number of channels, resulting in suboptimal performance.
In this work, we introduce a simple yet effective framework for long-term time series forecasting, AverageTime, which employs mixed-channel embedding, parameter-independent prediction, and prediction averaging before and after embedding to model time series data. Specifically, we first embed the channels of the input data with MLP and Transformer encoders to generate data with the same dimensionality as the original input. This process can be repeated several times to further extract features. Subsequently, we employ parameter-independent MLP layers for channel-wise prediction. Additionally, we introduce a clustering strategy to predict channels with high correlation together, thereby accelerating the process by shifting from channel-wise to group-wise predictions. The prediction results from the embedded channel vectors are first averaged and directly added to the predictions from the original data.
Experimental results indicate that AverageTime surpasses the performance of the current leading transformer architecture while maintaining efficiency comparable to lightweight linear models. The primary contributions of AverageTime can be summarized as follows.
@@
  • We introduce an averaging framework for long time series to capture inter-channel correlations. Specifically, we average the predictions of the original sequence with those obtained after channel embedding, achieving favorable performance in long-term time series forecasting.
@@
  • We develop a mixed-channel embedding layer that combines the distinct characteristics of Linear and Transformer layers to enhance the model’s ability to capture channel dependencies.
@@ 2 Problem Statement
In multivariate time series analysis, the objective is to solve the following problem: Given an input time series $X$ of dimensions $C\times T_{1}$, where $C$ is the number of channels and $T_{1}$ is the length of the time series. The future time period to be forecast is set to $T_{2}$. Therefore, a deep learning architecture is required to map the input time series $X$ to the predicted output $Y$ for $T_{2}$ future steps. This can be formalized as:
@@
where $f$ is a mapping function defined by a deep learning model that predicts $Y$ based on historical observations $X$. The function $f$ can be implemented using various architectures, including CNN, MLP, transformers, and their variants. Several methods have been developed to capture time series characteristics and enhance predictions, primarily categorized as follows.
@@ 2.1 Transformer Models
The Transformer architecture, an advanced method for long-term time series forecasting, excels at capturing dependencies within sequences and channels through the attention mechanism. The attention mechanism allows the model to focus on different parts of the input sequence when generating the output, efficiently handling sequences of varying lengths. The attention mechanism(Vaswani et al., 2017) is defined as:
@@
here, $Q$, $K$, and $V$ are feature matrices from the input data, and $d_{k}$ is the dimensionality of the keys. Recent research on transformer architectures has introduced several effective methods. PatchTST (Nie et al., 2023) divides the sequence into segments and applies attention, capturing relationships between different parts. iTransformer (Liu et al., 2024), embeds the channels and applies the attention mechanism to capture the complex dependency among channels, achieving significant performance improvements, especially in scenarios with a large number of channels. Additionally, methods like Informer (Zhou et al., 2021) address complexities in time series applications of attention, with these models demonstrating promising results.
@@ 2.2 MLP Models
In recent years, MLPs have become essential for multivariate long-term time series forecasting. With their straightforward architecture and strong predictive capabilities, MLPs play a pivotal role in multivariate time series forecasting. TiDE(Das et al., 2023) introduces a residual architecture into its linear layers to enhance the encoding of historical time series data and relevant covariates. The encoded information is then decoded with future covariates taken into account to generate forecasts. The dense encoder and decoder used in TiDE are shown as follows:
@@
where $M_{\phi}$ represents weight matrices, while $c_{\phi}$ denotes bias terms associated with the input time series. Through this approach, TiDE enhances the accuracy and robustness of predictions by leveraging both past data patterns and anticipated future conditions, thereby addressing the complexities inherent in time series forecasting. SparseTSF(Lin et al., 2024a) employs a cross-period modeling approach, wherein input data are first downsampled before being mapped through linear layers. The mapped data is then upsampled to generate the predicted time series output. Utilizing a linear model with fewer than 1k parameters, this approach achieves commendable forecasting performance. Such methodologies allow linear models to tackle time series forecasting with a highly efficient architecture, promoting their application in multivariate long-term forecasting tasks.
@@ 2.3 CNN Models
CNN-based time series analysis primarily relies on convolution operations to capture dependencies across both sequences and channels. MICN(Wang et al., 2022) achieves this through a multi-scale branch structure designed to model different underlying patterns separately, each branch extracts local features via downsampled convolution and captures global correlations through isometric convolution. Moreover, MICN exhibits linear complexity with respect to sequence length, making it more efficient than other models when dealing with long sequences. TimesNet(Wu et al., 2023) introduces the TimesBlock, a module that transforms 1D time series into 2D tensors. This transformation enables the modeling of complex temporal variations through efficient 2D convolution kernels, thereby effectively capturing both intra-series periodicities and inter-period variations. Furthermore, this approach is applicable to a variety of tasks within time series analysis, facilitating broader applications in time series forecasting and anomaly detection.
@@ 3.1 Overview
In this section, we first introduce the principal architecture of AverageTime, which is composed of MLPs for prediction, along with transformer encoders and MLPs for channel embedding. To capture the correlations among channels, we design a simple yet effective framework for long-term time series forecasting, as illustrated in Figure 1.
For the input time series, this approach aims to model channel dependencies using a combination of Transformer-based and MLP-based channel embedding layers, all the channel layers are shared across all time steps. As shown in the top of the figure, we first transpose the time series, apply multiple Transformer encoders and MLPs to capture the channel dependencies, the embedded time series is then transposed back, after which both the transposed time series and the original time series are processed using parameter-independent MLP layers. Finally, the outputs are averaged to obtain the final prediction. Additionally, we enhance the prediction process using a clustering method. Specifically, during prediction, a channel clustering approach is employed to ensure the channels with high correlations share the same MLP parameters. This transformation shifts channel-wise predictions to group-wise predictions, thereby improving model efficiency while maintaining predictive performance.
Figure 1: The overall architecture of the AverageTime model, which employs MLPs for predictions and incorporates mixed-channel embedding layer with shared parameters to capture inter-channel correlations.
Our model effectively leverages information both across channels and sequences through the use of a parameter-shared mixed-channel embedding layer and parameter-independent MLP prediction layers. It also incorporates the Revin regularization technique(Kim et al., 2021) to mitigate distributional shifts between sequences.
The core framework of our approach involves an averaging operation between the embedded results and the original predicted results. Averaging after prediction effectively decouples the channel and sequence capture processes, reducing the interaction between these two processes and thus improving the model’s performance.
@@ 3.2 LightAverageTime
For a long-term time series forecasting task, assuming we have an input time series $X^{C\cdot T}$, to calculate the correlation between different channels and ensure fairness, we first split the time series into ${X_{train}}^{C\cdot T_{train}}$, ${X_{test}}^{C\cdot T_{test}}$, ${X_{val}}^{C\cdot T_{val}}$, then we calculate spearman correlation on the training dataset, which is:
@@
where $r_{s}$ stands for the rank correlation, $d_{i}$ stands for the rank difference for different channels, $n$ stands for the number of channels in the train dataset. Through the application of spearman correlation, we can have an estimate of the similarity among channels $G^{C\cdot C}$, where $G^{i\cdot j}$ stands for the spearman correlation between channel $i$ and channel $j$, we then apply a method to filter the channels with low correlation with all other channels, in a manner similar to the behavior of an activation function, which is:
[…]
where threshold is a hyperparameter that we define to determine how a channel differs from all other channels. During the main experiment, the threshold is uniformly set to 0.8. After that, we use the Label Propagation Algorithm(Zhou et al., 2003) to group the channels. We first initialize a label for each channel $L_{i}$, where $i=1,2,3,\dots,C$. Then, we iterate over the labels, updating each label by taking the majority of the neighbor labels. Specifically, for the $t_{th}$ iteration, the label for the $i_{th}$ channel is:
$L_{i}^{(t)}=\arg\max_{L_{j}\in\{L_{1},L_{2},\dots,L_{n}\}}\left(\sum_{j\in\mathcal{N}(i)}\delta(L_{j},L_{i})\right),$ (6)
@@
[…]

### v4

@@ Abstract
Multivariate long-term time series forecasting aims to predict future sequences by utilizing historical observations, with a core focus on modeling intra-sequence and cross-channel dependencies. Numerous studies have developed diverse architectures to capture these patterns, achieving significant improvements in forecasting accuracy. Among them, iTransformer, a representative method for channel information extraction, leverages the Transformer architecture to model channel-wise dependencies, thereby facilitating sequence transformation for enhanced forecasting performance. Building upon iTransformer’s channel extraction concept, we propose AverageTime—a simple, efficient, and scalable forecasting model. Beyond iTransformer, AverageTime retains the original sequence information and reframes channel extraction as a stackable and extensible architecture. This allows the model to generate multiple novel sequences through various structural mechanisms, rather than being limited to transforming the original input. Moreover, the newly extracted sequences are not restricted to channel processing; other techniques such as series decomposition can also be incorporated to enhance predictive accuracy. Additionally, we introduce a channel clustering technique into AverageTime, which substantially improves training and inference efficiency with negligible performance loss. Experiments on real-world datasets demonstrate that with only two straightforward averaging operations—applied to both the extracted sequences and the original series—AverageTime surpasses state-of-the-art models in forecasting performance while maintaining near-linear complexity. This work offers a new perspective on time series forecasting: enriching sequence information through extraction and fusion. The source code is available at https://github.com/ UniqueoneZ/AverageTime.
keywords:
Long time series forecasting, Average, Cluster, Channel embedding
[…]
[label1]organization=School of Mathematics and Statistics, Shandong University, addressline=180 Wenhua West Road, city=Weihai, postcode=264209, state=Shandong, country=China
[…]
[label2]organization=Department of Mathematics, Harbin Institute of Technology (Weihai), addressline=2 Wenhua West Road, city=Weihai, postcode=264209, state=Shandong, country=China
[…]
[label3]organization=Research Center for Data Hub and Security, Zhejiang Lab, addressline=No. 1369 West Wenyi Road, city=Hangzhou, postcode=311121, state=Zhejiang, country=China
@@ 1 Introduction
Long-term time series prediction involves forecasting future trends over extended periods based on historical changes. This approach is crucial in various fields such as weather forecasting [11], traffic prediction [10], and power demand estimation [1]. The extended forecast horizon and the complex correlations across sequences and channels pose significant modeling challenges. Traditional methods often fall short of capturing sequence and inter-channel relationships. Consequently, the primary methodologies in this field have shifted to deep learning models.
@@
The core issue in long-term time series analysis is extracting dependencies within sequences and channels, which are crucial for improving model performance and robustness in multi-channel prediction. Various methods have been developed to capture these correlations from time series data. Widely employed techniques include Transformer-based models [18, 20, 24, 23, 12, 9, 16], which apply attention mechanisms to capture correlations both within sequences and across channels; model with CNN as the core architecture [14, 17], which use 1D or multidimensional convolutions to capture these dependencies; and structures based on Multilayer Perceptrons[19, 2, 8, 7, 15, 6], such as DLinear, which decompose sequences and apply multiple linear layers to predict future series. Among these, the iTransformer [9], which captures channel dependencies using channel embeddings and an attention structure, has achieved significant performance improvements on benchmark datasets, especially in datasets with a large number of channels. However, as the importance of channel correlations has been increasingly recognized [21], many models struggle to model the complex inter-channel correlations accurately, leading to a decline in model performance. This inadequacy becomes particularly evident in scenarios with a large number of channels, resulting in suboptimal performance.
In this work, we present AverageTime, a simple yet effective framework for multivariate long-term time series forecasting. It improves forecasting accuracy by fusing the original input series with complementary sequences that capture channel information derived from multiple architectural branches. Specifically, we apply multiple MLPs and Transformer Encoders along the channel dimension to capture diverse channel interactions, thereby generating several new sequences. These sequences are first averaged to perform initial information fusion. Subsequently, predictions from the fused sequence and the original sequence are averaged again, completing a second stage of information fusion. Additionally, we introduce a clustering strategy to predict channels with high correlation together, thereby accelerating the process by shifting from channel-wise to group-wise predictions. The prediction results from the embedded channel vectors are first averaged and directly added to the predictions from the original data. Experimental results indicate that AverageTime shows better performance than the current leading transformer architecture while maintaining efficiency comparable to lightweight linear models. The primary contributions of AverageTime can be summarized as follows.
  • 1.
@@
    We introduce an averaging framework for long time series to capture inter-channel information. Specifically, we average the predictions of the original sequence with those obtained after channel extraction, achieving favorable performance in long-term time series forecasting.
  • 2.
@@
  • 3.
    We design a scalable information fusion method that can extend channel-wise feature extraction to multiple architectures and sequences, while also allowing the introduction of additional information to enhance model effectiveness.
@@ 2 Problem Statement
In multivariate time series analysis, we aim to predict future values $Y\in\mathbb{R}^{C\times T_{2}}$ based on an input time series $X\in\mathbb{R}^{C\times T_{1}}$, by learning a function $f:\mathbb{R}^{C\times T_{1}}\rightarrow\mathbb{R}^{C\times T_{2}}$ , where $C$ is the number of channels and $T_{1}$ is the length of the time series. The future time period to be forecast is set to $T_{2}$. Therefore, a deep learning architecture is required to map the input time series $X$ to the predicted output $Y$ for $T_{2}$ future steps. This can be formalized as:
@@
where $f$ is a function defined by a deep learning model that predicts $Y$ based on historical observations $X$. The function $f$ can be implemented using various architectures, including CNN, MLP, transformers, and their variants. Several methods have been developed to capture time series characteristics and enhance predictions, primarily categorized as follows.
@@ 2.1 Transformer Models
The Transformer architecture, an advanced method for long-term time series forecasting, excels at capturing dependencies within sequences and channels through the attention mechanism. The attention mechanism allows the model to focus on different parts of the input sequence when generating the output, efficiently handling sequences of varying lengths. The attention mechanism[13] is defined as:
@@
here, $Q$, $K$, and $V$ are feature matrices from the input data, and $d_{k}$ is the dimensionality of the keys. Recent research on transformer architectures has introduced several effective methods. PatchTST [12] divides the sequence into segments and applies attention, capturing relationships between different parts. iTransformer [9], embeds the channels and applies the attention mechanism to capture the complex dependency among channels, achieving significant performance improvements, especially in scenarios with a large number of channels. Additionally, methods like Informer [23] address complexities in time series applications of attention, with these models demonstrating promising results.
@@ 2.2 MLP Models
In recent years, MLPs have become essential for multivariate long-term time series forecasting. With their straightforward architecture and strong predictive capabilities, MLPs play a pivotal role in multivariate time series forecasting. TiDE[3] introduces a residual architecture into its linear layers to enhance the encoding of historical time series data and relevant covariates. The encoded information is then decoded with future covariates taken into account to generate forecasts. The dense encoder and decoder used in TiDE are shown as follows:
@@
where $M_{\phi}$ represents weight matrices, while $c_{\phi}$ denotes bias terms associated with the input time series. Through this approach, TiDE enhances the accuracy and robustness of predictions by leveraging both past data patterns and anticipated future conditions, thereby addressing the complexities inherent in time series forecasting. SparseTSF[7] employs a cross-period modeling approach, wherein input data are first downsampled before being mapped through linear layers. The mapped data is then upsampled to generate the predicted time series output. Utilizing a linear model with fewer than 1k parameters, this approach achieves commendable forecasting performance. Such methodologies allow linear models to tackle time series forecasting with a highly efficient architecture, promoting their application in multivariate long-term forecasting tasks.
@@ 2.3 CNN Models
CNN-based time series analysis primarily relies on convolution operations to capture dependencies across both sequences and channels. MICN[14] achieves this through a multi-scale branch structure designed to model different underlying patterns separately, each branch extracts local features via downsampled convolution and captures global correlations through isometric convolution. Moreover, MICN exhibits linear complexity with respect to sequence length, making it more efficient than other models when dealing with long sequences. TimesNet[17] introduces the TimesBlock, a module that transforms 1D time series into 2D tensors. This transformation enables the modeling of complex temporal variations through efficient 2D convolution kernels, thereby effectively capturing both intra-series periodicities and inter-period variations. Furthermore, this approach is applicable to a variety of tasks within time series analysis, facilitating broader applications in time series forecasting and anomaly detection.
@@ 3.1 Overview
In this section, we introduce the proposed AverageTime method, which utilizes MLPs for sequence forecasting and employs transformer encoder and MLPs to extract inter-channel information to capture channel-wise dependencies; these dependencies are then fused to produce the final prediction. The framework is shown in Figure 1.
For the input time series, we first extract inter-channel dependencies using Transformer-based and MLP-based channel modules to generate new time series. As illustrated in the upper part of the figure, we initially transpose the time series and employ Transformer encoders along with MLPs to capture inter-channel dependencies, yielding a time series enriched with more channel information. This time series is then transposed back. Subsequently, we utilize parameter-independent MLP layers to predict all channels for each sequence individually, followed by fusing the predictions through averaging. Furthermore, to enhance the prediction process, we introduce a clustering strategy. Specifically, during prediction, highly correlated channels are assigned to MLP layers with shared parameters, improving prediction performance while minimizing interference. This shifts the approach from per-channel to per-channel-group prediction, significantly boosting inference efficiency.
Figure 1: The overall architecture of AverageTime, which employs MLPs and Transformer encoders to extract inter-channel information and form new time series. The final prediction is obtained by averaging the forecasts from the original sequence and the channel-augmented sequence.
Our model effectively leverages information both across channels and sequences. It also incorporates the Revin regularization technique[5] to mitigate distributional shifts between sequences.
The core of our framework lies in extracting new time series through channel interaction, which are then averaged with the original time series. This process effectively decouples the modeling of channel-wise and temporal dependencies. The information introduced by the newly extracted time series further enhances the prediction accuracy. Our model is designed to validate the feasibility of constructing and integrating such sequences. The fusion strategy may also be extended to more sophisticated mechanisms, such as gating operations, to further improve model performance.
@@ 3.2 LightAverageTime
For a long-term time series forecasting task, assuming we have an input time series $X^{C\times T}$, to calculate the correlation between different channels and ensure fairness, we first split the time series into ${X_{train}}^{C\times T_{train}}$, ${X_{test}}^{C\times T_{test}}$, ${X_{val}}^{C\times T_{val}}$, then we calculate the Spearman correlation on the training dataset. It should be noted that other correlation coefficients can also be adopted; however, their impact on the results is generally minor. The equation of the Spearman correlation is given as follows:
@@
where $r_{s}$ stands for the rank correlation, $d_{i}$ stands for the rank difference for different channels, $n$ stands for the number of channels in the train dataset. Through the application of Spearman correlation, we can have an estimate of the similarity among channels $G^{C\times C}$, where $G^{i\times j}$ stands for the Spearman correlation between channel $i$ and channel $j$, we then apply a method to filter the channels with low correlation with all other channels, in a manner similar to the behavior of an activation function, which is:
$G^{i\times j}=\begin{cases}1,&\text{if }G^{i\times j}>T\\ 0,&\text{if }G^{i\times j}<T\end{cases}$ (5)
where T is a threshold hyperparameter that we define to determine how a channel differs from all other channels. During the main experiment, the $T$ is uniformly set to 0.8. After that, we use the Label Propagation Algorithm[22] to group the channels. We first initialize a label for each channel $L_{i}$, where $i=1,2,3,\dots,C$. Then, we iterate over the labels, updating each label by taking the majority of the neighbor labels. Specifically, for the $t^{th}$ iteration, the label for the $i^{th}$ channel is:
$L_{i}^{(t)}=\arg\max_{L_{j}\in\{L_{1},L_{2},\dots,L_{C}\}}\left(\sum_{j\in\mathcal{N}(i)}\delta(L_{j},L_{i})\right),$ (6)
@@
$\delta(L_{j},L_{i})=\begin{cases}1,&\text{if }L_{i}=L_{j}\\ 0,&\text{if }L_{i}\neq L_{j}\end{cases}$ (7)
@@
3.3 FVMD
The core of AverageTime is to incorporate more information and fuse it through simple averaging. Previously, we extracted channel information to form new sequences, which were then fused with the original sequences. These new sequences are not limited to those derived from channel extraction; they can be provided to the model from multiple perspectives. We further integrate information from the sequence decomposition method VMD[4] as a branch of the model. The optimization objective of VMD is to iteratively decompose the original signal $f$ into $K$ mode components $u_{k}(t)$ with limited bandwidth, ensuring that the sum of all modes equals the original signal and minimizing the total estimated bandwidth of the modes. The objective function is:
$\min_{\{u_{k}\},\{\omega_{k}\}}\left\{\sum_{k=1}^{K}\left\|\partial_{t}\left[\left(\delta(t)+\frac{j}{\pi t}\right)*u_{k}(t)\right]e^{-j\omega_{k}t}\right\|_{2}^{2}\right\}\quad\text{s.t.}\quad\sum_{k=1}^{K}u_{k}(t)=f(t)$
where $\delta(t)$ denotes the Dirac delta function, $j$ is the imaginary unit, and * represents the convolution operator. The operator $\left(\delta(t)+\frac{j}{\pi t}\right)*(\cdot)$ corresponds to the construction of the analytic signal via the Hilbert transform. The exponential term $\mathrm{e}^{-j\omega_{k}t}$ shifts the spectrum of the $k$-th mode to the baseband, while the time derivative $\partial_{t}$ is used to measure the bandwidth of each mode. By introducing a quadratic penalty term and a Lagrange multiplier, an augmented Lagrangian function is constructed and optimized via the Alternating Direction Method of Multipliers (ADMM).
The core of long-term time series forecasting lies in generalization capability. To this end, we introduce a modification to VMD by incorporating a penalty term into its objective function to constrain the center frequencies of the decomposed modes, named FVMD. The penalty term is given by:
$\Phi(\{w_{k}\})=\eta\sum_{k=2}^{K}\frac{1}{w_{k}-w_{k-1}}$
where $\eta>0$ is a regularization parameter that controls the strength of the frequency separation constraint. Following the optimization approach of VMD, we employ the ADMM for optimization. Initialize $\omega_{k}^{(0)}$, $\hat{u}_{k}^{(0)}(\omega)$, and $\hat{\lambda}^{(0)}(\omega)$. For each mode $k$, update the mode function:
$\hat{u}_{k}^{(n+1)}(\omega)=\frac{\hat{f}(\omega)-\sum_{i\neq k}\hat{u}_{i}^{(n)}(\omega)+\frac{\hat{\lambda}^{(n)}(\omega)}{\alpha}}{1+\frac{2}{\alpha}(\omega-\omega_{k}^{(n)})^{2}}$
Update the center frequency:
$\omega_{k}^{(n+1)}=\frac{B}{A}+\frac{\eta}{2A}\left(\frac{1}{(\omega_{k}^{(n)}-\omega_{k-1}^{(n)})^{2}}+\frac{1}{(\omega_{k+1}^{(n)}-\omega_{k}^{(n)})^{2}}\right)$
Update the Lagrange multiplier:
$\hat{\lambda}^{(n+1)}(\omega)=\hat{\lambda}^{(n)}(\omega)+\tau\left(\hat{f}(\omega)-\sum_{k}\hat{u}_{k}^{(n+1)}(\omega)\right)$
FVMD aims to provide additional information for the AverageTime framework, primarily to validate the effectiveness of multi-information fusion, rather than to highlight its comparison with the standard VMD method.